# Deception Against Data-Driven Linear-Quadratic Control

### v1

@@ 1 Introduction
To obtain the defender’s deceptive input, we cast its objective as a constrained optimization problem balancing two objectives: i) the desire to steer the adversary’s learned policy towards an a priori chosen benign gain; and ii) the desire to use as small of deceptive feedback as possible to maintain the nominal closed-loop performance and stability. We mathematically specify the conditions under which the defender’s optimization is well-defined and characterize its solution as the root of a coupled algebraic Riccati equation and a Lyapunov equation. Subsequently, since these coupled equations are difficult to analytically solve, we tackle them numerically with a successive over-relaxation iterative algorithm. In addition, we prove this algorithm’s convergence under certain conditions. We finally note that the proposed deception design also applies to the easier problem of deceiving minimizing data-driven linear-quadratic regulators.
@@ 2 Preliminaries
Notation: We denote as $\mathbb{R}$ the set of real numbers, and as $\mathbb{N}$ the set of natural numbers including zero. The vector $e_{i}\in\mathbb{R}^{n}$ denotes a unit vector with a unity located in the $i$-th row. For any matrix $X$, $\left\lVert X\right\rVert_{\mathrm{F}}$ denotes its Frobenius norm, whereas $\mathrm{Ran}(X)$ denotes its range. In addition, $[X]_{ij}$ denotes the entry in its $i$-th row and $j$-th column. Moreover, we follow the matrix differentiation convention $\left[\frac{\mathrm{d}}{\mathrm{d}X}\right]_{ij}=\frac{\mathrm{d}}{\mathrm{d}[X]_{ij}}$. For a square matrix $X$, $\alpha(X)$ denotes its spectral abscissa, and $\textrm{tr}(X)$ its trace. For a symmetric matrix $X$, $\lambda_{\mathrm{min}}(X)$ and $\lambda_{\mathrm{max}}(X)$ denote its minimum and maximum eigenvalue, respectively. We denote $X\succ 0$ (respectively, $X\succeq 0)$ if $X$ is positive definite (respectively, positive semidefinite). We denote as $\textrm{vec}(X)$ the vectorization of $X$. We use $I_{n}$ to denote the identity matrix of order $n$ (we omit the subscript when the order is obvious). Finally, we denote the Kronecker delta function as $\delta_{i,j}$, so that $\delta_{i,j}=1$ if $i=j$, and $\delta_{i,j}=0$ else.
@@
Consider the following algebraic Riccati equation:
@@ 3.2 The Adversary’s Model and Objectives
where $Q,R\succ 0$ are weighting matrices and $K$ is the adversary’s attack gain that is to be optimized. Provided that $\lambda_{\textrm{min}}(R)$ is above a threshold, it then follows from standard linear systems theory [26] that the solution to this problem is given by
@@
    The adversary does not know the values of the entries of the system’s matrices $A,B_{a}$.
@@ 3.3 The Defender’s Model and Objectives
We assume that the defender of the system is aware that a potential adversary might be gathering information from the system with the aim of synthesizing the optimal attack gain $K^{\star}$ in (5). In that respect, the defender opts to spoof the system’s dynamics during this window, and deceive the adversary into learning a prespecified gain different from $K^{\star}$. Specifically, during this window, the system’s dynamics take the form
@@ 4.1 Deception Design using a Distance-Based Metric
where $\bar{K}$ is as in Problem 1, $\Gamma\succ 0$, and $\mathrm{tr}(\Lambda^{\mathrm{T}}\Gamma\Lambda)$ is a regulation term. This regulation term is necessary to ensure that the minimum exists, but also to penalize how large the deception gain $\Lambda$ can be to minimize performance loss and maintain stability [^1] This is for the case where $\bar{K}$ is not stabilizing. For the case where $\bar{K}=0$, regularization is not needed; see Section 6.1. . As an alternative, the distance of the target gain $\bar{K}$ from the original gain $K^{\star}$ should be sufficiently small, assuming $K^{\star}$ exists. We provide an explicit – though conservative – condition on the distance $\left\lVert K^{\star}-\bar{K}\right\rVert_{\mathrm{F}}$ as well as the regulation weight $\Gamma$, for ensuring closed-loop stability and the feasibility of (11), in the following Lemma. The Lemma’s proof, as well as the proofs of all technical results, are postponed to the Appendix.
@@ 5 Convergence Analysis
The practical interpretation of Lemma 3 is that, by continuity, $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists not only everywhere in $\mathcal{S}_{0}$ but also in a neighborhood around it. Therefore, the negated gradient $-\frac{\mathrm{d}\tilde{J}(\Lambda)}{\mathrm{d}\Lambda}$ points in a direction towards which $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists, and hence this direction is in the interior of $\mathcal{S}_{0}$. Accordingly, Algorithm 1 will always remain within a set where both equations (15) and (16) admit unique solutions and will not terminate prematurely. This effectively deals with one of the two issues raised in the preceding paragraphs.
@@ 6.1 On the Condition of Lemma 1
Consider a scalar system where the closed-loop state matrix is $A=-1$, the input matrices are $B_{u}=B_{a}=1$, and the state cost matrix is $Q=1$. Then, then ARE (6) reduces to the scalar quadratic equation
@@
When $R>1$ solution, the solutions to this equation are given by $P_{1,2}=R\pm\sqrt{R-1}\sqrt{R}$, and hence the minimal positive definite solution is $P=R-\sqrt{R-1}\sqrt{R}$. Therefore, from (5), the optimal attack gain is $K^{\star}=1-\sqrt{1-R^{-1}}$. On the other hand, when $10<R<$ 1, (26) admits no solutions, the value of (4) is unbounded, and the optimal attack $K^{\star}$ destabilizing. Yet, the defender can deceive the attacker into a milder attack that renders the closed-loop stable, while bypassing the rationale of Lemma 1.
@@ 7 Simulations
We consider a linearized model of the ADMIRE benchmark aircraft [32], whose open-loop state and input matrices $A_{o},B_{u}$ are given in the Appendix. Using these, we define the closed-loop state matrix of the aircraft as $A=A_{o}+B_{u}F$, where the gain $F$ originates from a linear-quadratic regulation problem with identity weighting matrices.
@@ 7.1 Case 1: Deception against a Performance-Degrading Attack
Figure 3: Case 1: The evolution of the cost (11) during Algorithm 1.
@@
The evolutions of the deception gain sequence $\Lambda^{i}$ generated by Algorithm 1, as well as of the deception cost (11), are shown in Figures 2-3. From Figure 2, we clearly notice that Algorithm 1 reaches a stationary point of the deception problem (11) after about 3500 iterations, which validates Theorem 2. In addition, Figure 3 shows that the deception cost (11) is decreasing at each iteration of Algorithm 1. This is in line with the observation made in Lemma 2, regarding the Algorithm’s equivalence to a first-order gradient method. It indicates that, at each iteration $i$, the gain $\Lambda^{i}$ the defender computes becomes more and more efficient in deceiving the adversary. Table 1 additionally shows the ratio between the optimal attack $K^{\star}$ the attacker would have learned in the absence of deception, and the attack $K_{u}(\Lambda^{\star})$ they were forced to learn instead. Clearly, the latter has been suppressed substantially and is thus more benign in disturbing closed-loop performance. Note that $K_{u}(\Lambda^{\star})$ does not become exactly zero because of the regularization term that penalizes large values of $\Lambda$ in (11).
@@
Table 1: Ratio $[K_{u}(\Lambda^{\star})]_{ij}/[K^{\star}]_{ij}$ between optimal and deceived attack gains.
Col. $1$ Col. $2$ Col. $3$ Col. $4$ Col. $5$
Row $1$ $0.0827$ $0.9429$ $0.0545$ $0.2748$ $0.2439$
Row $2$ $0.1581$ $0.0046$ $0.2179$ $0.1103$ $0.2189$
@@
Table 2: Ratio $[\hat{K}_{u}(\Lambda^{\star})]_{ij}/[\hat{K}^{\star}]_{ij}$ between optimal and deceived attack gains, under incorrectly assumed weighing matrices for the adversary.
Col. $1$ Col. $2$ Col. $3$ Col. $4$ Col. $5$
Row $1$ $0.0655$ $-0.0829$ $0.0355$ $0.2159$ $0.2531$
Row $2$ $0.1190$ $0.0014$ $0.1761$ $0.0694$ $0.1366$
@@
Next, we showcase the robustness of the derived deception gain when the adversary’s weighing matrices are different than those assumed. Specifically, we take the deception gain we previously derived under $Q=I_{5}$ and $R=0.6I_{2}$, and show that even if the adversary’s weighing matrices are equal to $\hat{Q}=5I_{6}\neq Q$ and $\hat{R}=2I_{2}\neq R$, then the learned adversarial attack is still suppressed. We show the ratio between the learned adversarial attack $\hat{K}_{u}(\Lambda^{\star})$ and the nominal attack $\hat{K}^{\star}$ under $\hat{Q},~{}\hat{R}$ in Table 2, validating this conjecture along with the analysis we carried out in Section 6.2.
7.2 Case 2: Deception against a Destabilizing Attack
In Case 2, we consider the same setup for the adversary, but where $R=0.1I_{2}$ in their optimization problem (4). In this case, the optimal attack $K^{\star}$ is destabilizing. To find a deception gain that will force the adversary to learn a more benign attack instead, we re-employ the same scheme as in Case 1 for the defender. However, following the discussion in Section 6, we select $\Lambda^{0}$ in Algorithm $1$ so that it makes the eigenvalues of $A+B_{u}\Lambda^{0}$ sufficiently negative; in particular, equal to $-100$.
The evolution of the deception gain sequence $\Lambda^{i}$ and the deception cost (11) during Algorithm 1 is shown in Figures 4-5. While we still observe a decreasing deception cost and a convergent deception gain, the number of iterations needed to achieve the same level of tolerance increased significantly. This was due to the conservative initialization of $\Lambda^{0}$ very far from the optimal deception gain $\Lambda^{\star}$. Meanwhile, the eigenvalues of the state matrix in closed-loop with the deceptive attack gain are equal to $-17.6911,-7.7253,-2.5628+0.4642i,-2.5628-0.4642i,-1.6568$. This means that the learned optimal attack under deception is not destabilizing, as opposed to the nominal optimal attack $K^{\star}$.

### v2

@@ 1 Introduction
From a control-theoretic perspective, deception in dynamical systems can be formulated within a broader optimal control and game-theoretic framework. In such settings, one agent strategically perturbs information or system inputs in order to influence the policy computed or learned by one or more other agents [6]. This leads naturally to bilevel optimal control problems in which one decision maker anticipates and shapes the optimization or learning process of another.
@@
To obtain the defender’s deceptive input, we cast its objective as a constrained optimization problem balancing two objectives: i) the desire to steer the adversary’s learned policy towards an a priori chosen benign gain; and ii) the desire to use as small of deceptive feedback as possible to maintain the nominal closed-loop characteristics and stability. We mathematically specify the conditions under which the defender’s optimization is well-defined and characterize its solution as the root of a coupled algebraic Riccati equation and a Lyapunov equation. Subsequently, since these coupled equations are difficult to analytically solve, we tackle them numerically with a successive over-relaxation iterative algorithm. In addition, we prove this algorithm’s convergence under certain conditions. We finally note that the proposed deception design also applies to the easier problem of deceiving minimizing data-driven linear-quadratic regulators.
@@
The contributions of the paper are summarized as follows.
  1. 1.
    We consider deception in a setting where a defender, who has knowledge of the system’s matrices, injects an additional state-feedback input to strategically influence the learning dynamics of a data-driven attacker, with the goal of steering it toward a suboptimal attack policy. Unlike classical LQR game formulations [27, 28, 29], where all players have model access and the equilibrium is uniquely determined by the problem parameters, our setting exploits the fact that the adversary must learn its policy from data, opening the door to shaping what it converges to. This also distinguishes our approach from moving target defense [24, 25, 26], which does not explicitly target the adversary’s learned policy, and from deception in Nash equilibrium seeking [11], which considers static game settings.
  2. 2.
    We characterize the optimal deception gain as the solution of a coupled algebraic Riccati and Lyapunov equation, and develop a block successive over-relaxation algorithm with provable convergence guarantees. This provides theoretical foundations for this specific problem that are absent from related data poisoning approaches, such as those on virtual reference feedback controllers [17, 18] and data predictive control [19], as well as from the preliminary version of this paper [30].
  3. 3.
    We analyze robustness to uncertainty in the adversary’s cost matrices and validate the framework on an aircraft model, demonstrating substantial attack suppression.
Structure: The remainder of the paper is organized as follows. Section 2 presents preliminaries, Section 3 formulates the deception problem, and Section 4 outlines the proposed solution within an algorithmic framework. Section 5 establishes convergence guarantees for the algorithm, while Section 6 discusses robustness to parametric uncertainty and relaxed assumptions. Finally, Section 7 presents simulation results, and Section 8 concludes the paper.
@@ 2 Preliminaries
Notation: We denote as $\mathbb{R}$ the set of real numbers, and as $\mathbb{N}$ the set of natural numbers including zero. The vector $e_{i}\in\mathbb{R}^{n}$ denotes a unit vector with a unity located in the $i$-th row. For any matrix $X$, $\left\lVert X\right\rVert_{\mathrm{F}}$ denotes its Frobenius norm, whereas $\mathrm{Ran}(X)$ denotes its range (i.e., its column space). In addition, $X]_{ij}$ denotes the entry in its $i$-th row and $j$-th column. Moreover, we follow the matrix differentiation convention $\left[\frac{\mathrm{d}}{\mathrm{d}X}\right]_{ij}=\frac{\mathrm{d}}{\mathrm{d}[X]_{ij}}$. For a square matrix $X$, $\alpha(X)$ denotes its spectral abscissa, and $\textrm{tr}(X)$ its trace. For a symmetric matrix $X$, $\lambda_{\mathrm{min}}(X)$ and $\lambda_{\mathrm{max}}(X)$ denote its minimum and maximum eigenvalue, respectively. We denote $X\succ 0$ (respectively, $X\succeq 0)$ if $X$ is positive definite (respectively, positive semidefinite). We denote as $\textrm{vec}(X)$ the vectorization of $X$. We use $I_{n}$ to denote the identity matrix of order $n$ (we omit the subscript when the order is obvious). We denote the Kronecker delta function as $\delta_{i,j}$, so that $\delta_{i,j}=1$ if $i=j$, and $\delta_{i,j}=0$ else.
@@
Consider the algebraic Riccati equation:
@@ 3.2 The Adversary’s Model and Objectives
where $Q,~R\succ 0$ are weighting matrices and $K$ is the adversary’s attack gain that is to be optimized. Provided that $\lambda_{\textrm{min}}(R)$ is above a threshold, it then follows from standard linear systems theory [27] that the solution to this problem is given by
@@
    The adversary does not know the values of the entries of the system matrix $A$. They may or may not know the values of the entries of the system matrix $B_{a}$.
@@ 3.3 The Defender’s Model and Objectives
We assume that the defender of the system is aware that a potential adversary might be gathering information from the system with the aim of synthesizing the optimal attack gain $K^{\star}$ in (5). In response, the defender may temporarily modify the system’s dynamics to induce the adversary to learn a prespecified gain different from $K^{\star}$. Specifically, during this window, the system’s dynamics take the form
@@ 4.1 Deception Design using a Distance-Based Metric
where $\bar{K}$ is as in Problem 1, $\Gamma\succ 0$, and $\mathrm{tr}(\Lambda^{\mathrm{T}}\Gamma\Lambda)$ is a regulation term. This regulation term is necessary to ensure that the minimum exists, but also to penalize how large the deception gain $\Lambda$ can be to minimize deviation from nominal closed-loop performance and maintain stability [^1] This is for the case where $\bar{K}$ is not stabilizing. For the case where $\bar{K}=0$, regularization is not needed; see Section 6.1. . As an alternative, the distance of the target gain $\bar{K}$ from the original gain $K^{\star}$ should be sufficiently small, assuming $K^{\star}$ exists. We provide an explicit – though conservative – condition on the distance $\left\lVert K^{\star}-\bar{K}\right\rVert_{\mathrm{F}}$ as well as the regulation weight $\Gamma$, for ensuring closed-loop stability and the feasibility of (11), in the following Lemma. The Lemma’s proof, as well as the proofs of all technical results, are postponed to the Appendix.
@@ 5 Convergence Analysis
The practical interpretation of Lemma 3 is that, by continuity, $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists not only everywhere in $\mathcal{S}_{0}$ but also in a neighborhood around it. Therefore, the negated gradient $\frac{\mathrm{d}\tilde{J}(\Lambda)}{\mathrm{d}\Lambda}$ points in a direction towards which $P_{u}(\Lambda)\in\mathbb{S}_{+}$ exists, and hence this direction is in the interior of $\mathcal{S}_{0}$. Accordingly, Algorithm 1 will always remain within a set where both equations (15) and (16) admit unique solutions and will not terminate prematurely. This effectively deals with one of the two issues raised in the preceding paragraphs.
@@ 6.1 On the Condition of Lemma 1
Consider a scalar system where the closed-loop state matrix is A=1A=-1, the input matrices are $B_{u}=B_{a}=1$, and the state cost matrix is $Q=1$. Then, then ARE (6) reduces to the scalar quadratic equation
@@
When $R>1$ solution, the solutions to this equation are given by $P_{1,2}=R\pm\sqrt{R-1}\sqrt{R}$, and hence the minimal positive definite solution is $P=R-\sqrt{R-1}\sqrt{R}$. Therefore, from (5), the optimal attack gain is $K^{\star}=1-\sqrt{1-R^{-1}}$. On the other hand, when $0<R<1$, (26) admits no solutions, the value of (4) is unbounded, and the optimal attack $K^{\star}$ destabilizing. Yet, the defender can deceive the attacker into a milder attack that renders the closed-loop stable, while bypassing the rationale of Lemma 1.
@@ 6.2 On the Assumption of Known $Q$ and $R$ Matrices
Remark 6.
Proposition 2 establishes that $\hat{J}(\Lambda^{\star})\leq\tilde{J}(\Lambda^{\star})+\epsilon$. The gap $\hat{J}(\Lambda^{\star})-\tilde{J}(\Lambda^{\star})$ depends continuously on the cost-matrix mismatches $\|\hat{Q}-Q\|_{\textrm{F}}$ and $\|\hat{R}-R\|_{\textrm{F}}$ and vanishes as these mismatches vanish. This follows from standard norm inequalities together with the continuous dependence of the stabilizing algebraic Riccati solution (and hence of the induced feedback gain) on its cost matrices.
@@ 7 Simulations
We consider a linearized model of the ADMIRE benchmark aircraft, whose open-loop state and input matrices $A_{o},B_{u}$ are given in [33]. Using these, we define the closed-loop state matrix of the aircraft as $A=A_{o}+B_{u}F$, where the gain $F$ originates from a linear-quadratic regulation problem with identity weighting matrices.
Figure 3: Case 1: The evolution of the cost (11) during Algorithm 1.
Figure 4: Impact of deception on the evolution of the learning algorithms presented in [34, 35]. The figure shows the distance of the learnt gain $K_{j}$ at each iteration $j$ of the learning algorithm, from the suboptimal gain $K_{u}(\Lambda^{\star})$.
@@ 7.1 Case 1: Deception against a Performance-Degrading Attack
The evolutions of the deception gain sequence $\Lambda^{i}$ generated by Algorithm 1, as well as of the deception cost (11), are shown in Figures 2-3. From Figure 2, we notice that Algorithm 1 reaches a stationary point of the deception problem (11) after about 3500 iterations, which validates Theorem 2. In addition, Figure 3 shows that the deception cost (11) is decreasing at each iteration of Algorithm 1. This is in line with the observation made in Lemma 2, regarding the Algorithm’s equivalence to a first-order gradient method. It indicates that, at each iteration $i$, the gain $\Lambda^{i}$ the defender computes becomes more and more efficient in deceiving the adversary.
@@
Table 1: Ratio $K_{u}(\Lambda^{\star})]_{ij}/[K^{\star}]_{ij}$ between optimal and deceived attack gains.
Col. 11 Col. 22 Col. 33 Col. 44 Col. 55
Row 11 $0.0827$ $0.9429$ $0.0545$ $0.2748$ $0.2439$
Row 22 $0.1581$ $0.0046$ $0.2179$ $0.1103$ $0.2189$
@@
Table 2: Ratio $\hat{K}_{u}(\Lambda^{\star})]_{ij}/[\hat{K}^{\star}]_{ij}$ between optimal and deceived attack gains, under incorrectly assumed weighing matrices for the adversary.
Col. 11 Col. 22 Col. 33 Col. 44 Col. 55
Row 11 $0.0655$ 0.0829-0.0829 $0.0355$ $0.2159$ $0.2531$
Row 22 $0.1190$ $0.0014$ $0.1761$ $0.0694$ $0.1366$
Table 3: State energy $\mathcal{E}_{x}=\int_{0}^{\infty}\left\lVert x(t)\right\rVert^{2}\,dt$ under nominal and deceived attacks for different adversarial input weights $R$. The value $\infty$ indicates closed-loop instability.
Case $R=0.2I_{2}$ $R=0.4I_{2}$ $R=0.6I_{2}$ $R=0.8I_{2}$
Nominal $\infty$ $0.022844$ $0.008229$ $0.008055$
Deceived $0.007948$ $0.007924$ $0.007914$ $0.007909$
Table 1 additionally shows the ratio between the optimal attack $K^{\star}$ the attacker would have learned in the absence of deception, and the attack $K_{u}(\Lambda^{\star})$ they were forced to learn instead. Clearly, the latter has been suppressed substantially and is thus more benign in disturbing closed-loop performance. Note that $K_{u}(\Lambda^{\star})$ does not become exactly zero because of the regularization term that penalizes large values of $\Lambda$ in (11).
To validate this prediction at the level of the attacker’s learning dynamics, Figure 4 illustrates how the deception gain $\Lambda^{\star}$ affects the evolution of two learning algorithms from the literature [34, 35]. One assumes unknown $A$ and $B_{a}$ matrices [35], whereas the other assumes unknown $A$ and known $B_{a}$ [34]. In both cases, we observe the learning process converges to the suboptimal attack $K_{u}(\Lambda^{\star})$ that the defender had predicted.
To further assess the impact at the level of closed-loop performance, Table 3 reports the resulting state energy $\mathcal{E}_{x}=\int_{0}^{\infty}\left\lVert x(t)\right\rVert^{2}dt$ under nominal and deceived attacks for different values of $R$. We observe that, for small values of $R$, the nominal attack may even destabilize the system (see next subsection), whereas the deceived attack consistently yields finite and smaller energy.
@@
Finally, we showcase the robustness of the derived deception gain when the adversary’s weighing matrices are different than those assumed. Specifically, we take the deception gain we previously derived under $Q=I_{5}$ and $R=0.6I_{2}$, and show that even if the adversary’s weighing matrices are equal to $\hat{Q}=5I_{6}\neq Q$ and $\hat{R}=2I_{2}\neq R$, then the learned adversarial attack is still suppressed. We show the ratio between the learned adversarial attack $\hat{K}_{u}(\Lambda^{\star})$ and the nominal attack $\hat{K}^{\star}$ under $\hat{Q},~\hat{R}$ in Table 2, validating this conjecture along with the analysis we carried out in Section 6.2.
@@
7.2 Case 2: Deception against a Destabilizing Attack
In Case 2, we consider the same setup for the adversary, but where $R=0.1I_{2}$ in their optimization problem (4). In this case, the optimal attack $K^{\star}$ is destabilizing. To find a deception gain that will force the adversary to learn a more benign attack instead, we re-employ the same scheme as in Case 1 for the defender. However, following the discussion in Section 6, we select $\Lambda^{0}$ in Algorithm $1$ so that it makes the eigenvalues of $A+B_{u}\Lambda^{0}$ sufficiently negative; in particular, equal to 100-100.
The evolution of the deception gain sequence $\Lambda^{i}$ and the deception cost (11) during Algorithm 1 is shown in Figures 5-6. While we still observe a decreasing deception cost and a convergent deception gain, the number of iterations needed to achieve the same level of tolerance increased significantly. This was due to the conservative initialization of $\Lambda^{0}$ very far from the optimal deception gain $\Lambda^{\star}$. Meanwhile, the eigenvalues of the state matrix in closed-loop with the deceptive attack gain are equal to 17.6911,7.7253,2.5628+0.4642i,2.56280.4642i,1.6568-17.6911,-7.7253,-2.5628+0.4642i,-2.5628-0.4642i,-1.6568. This means that the learned optimal attack under deception is not destabilizing, as opposed to the nominal optimal attack $K^{\star}$.